# Automatic Organization of Neural Modules for Enhanced Collaboration in Neural Networks

### v2

@@ 2 Related Works
To progress the existing tree-like structure for NNs. Yuan 2020ECCV recently provided a topological perspective, highlighting the benefits of dense connections offered through shortcuts in optimization 2015Highway 2018MobileNetV2. Furthermore, sparsity constraints have also been proven effective in optimizing learned structures across various applications 2015Highway chu2023learning suc101 suc102 suc103. In their approach, the structure of NNs is organized as a DAG, whereas we organize it as a more general graph structure.
@@
The fixed point of the implicitly hidden layer can also be served as a solution bai2019deep tsuchida2022equilibrium chu2023learning yang2022closer heaton2021feasibility zucchet2022beyond, as demonstrated in the subsequent works bai2020multiscale szekeres2022computation suc104. Departing from the infinite structure of implicitly hidden layers chu2023learning ling2023equivalence ding2023two yang2022closer liu2022online, we organize it into a general graph structure. Furthermore, for larger NNs, we introduce a novel regularization method to organize neural units into multiple independent neural modules. Compared with implicitly hidden layers, our method can improve the efficiency by parallel computing as well as the performance by reducing overfitting. Furthermore, it enhances the interpretability of implicitly hidden layers.
@@
Our process also involves compressing NNs. In recent years, various algorithms have been developed, including quantization suc1 suc2 suc3, low-rank approximation suc4 suc5 suc6, knowledge distillation suc7 suc8, and network pruning suc88, etc. The network pruning method we use in this work is weight pruning, which aims to eliminate weak connections. In addition, the method also involves filter pruning that removes entire redundant filters suc9. We aim to improve weight pruning for our framework using a method similar to suc10, which evaluates the gradient at the pruned model and applies parameters’ updates to the dense model. In our framework, this process coordinates with an elegant regularization to automatically allocate Neural Modules.
@@
OptNet integrates optimization quadratic problems for nodes within the same layer suc11 suc12. However, this approach introduces bias and lacks an interpretable structure, affecting overall interpretability. Furthermore, OptNet uses parameters derived from quadratic problems for backpropagation. These parameters, determined by nodes from the previous layer, are difficult to manage in complexity. This complexity not only increases the computational burden but also exacerbates overfitting problems. Using a differentiable function for these parameters adds further bias.
@@
Graph Neural Networks (GNNs) specially address the needs of geometric deep learning 2005A 2019Graph scarselli2008graph abadal2021computing. GNNs adapt their structure to the input graph, capturing complex dependencies yong2007document abadal2021computing 2017Protein 2019Graph. This adaptability enables them to predict properties of geometric data, which primarily deal with graphs, our innovation involves evolving the tree-like structure into a general graph format for NNs, expanding their applicability and potential.
@@ 3.1 Rationale for Introducing the Structure
Consider the traditional tree-like structure for an NN with $m$ layers. Let $N^{0}$ be the nodes of the first layer and the input variables feed into the first layer. Let $N^{m}$ be nodes for the last layer and they feed for the output variables. For any node in the other layers, $n^{k}_{i}\in N^{k},0<k<m$, it represents the intermediate variables. The parameters of the model are represented as the edges in the structure. Let $e^{k}_{i,j}\in E^{k},1\leq k\leq m$ be the the edge connecting $n^{k-1}_{i}$ and $n^{k}_{j}$. Then the traditional tree-like structure for NNs which is organized as an asynchronous hierarchical structure can be formalized as follows: $\mathcal{T}=\{N^{0},E^{1},N^{1},...E^{m},N^{m}\}$.
@@ 3.3 Forward Process
Existing numerical methods can effectively solve these above equations. In real-world applications, efficiency can be optimized. Note that, each variable is processed by an activation function $f$, making the transformation nonlinear. If we consider $\mathcal{G}$ as a Neural Module, we can introduce the following theorem.
@@ 3.4 Backward Process
$\begin{cases}\sum\limits_{j\neq 1}\nabla x_{j}\cdot f^{-1}(x_{j})\cdot w_{1,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{-1}(x^{m}_{j})\cdot w^{m}_{1,j}=\nabla x_{1}\\ \sum\limits_{j\neq 2}\nabla x_{j}\cdot f^{-1}(x_{j})\cdot w_{2,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{-1}(x^{m}_{j})\cdot w^{m}_{2,j}=\nabla x_{2}\\ ...\\ \sum\limits_{j\neq m}\nabla x_{j}\cdot f^{-1}(x_{j})\cdot w_{p,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{-1}(x^{m}_{j})\cdot w^{m}_{p,j}=\nabla x_{p}\\ \end{cases}$
@@
$\nabla W^{T}_{:,j}=\nabla x_{j}\cdot f^{-1}(x_{j})\cdot\mathcal{H}_{j}~{}.$ (2)
@@
$\nabla W^{m}=\nabla X^{mT}\cdot g^{-1}(X^{mT})\cdot f(X)~{}.$ (3)
@@ 40 Nodes 60 Nodes 80 Nodes 100 Nodes 200 Nodes 300 Nodes
NMs&NM 0.0709 0.2416 0.0724 0.0570 0.0631 0.0635
4.2 The Optimization of the Neural Modules
In this section, we explore the process of optimizing the structure of Neural Modules using $NM$ regularization. Since every structure is a subgraph of a fully connected graph, the initial general graph serves as the search space for our model. As discussed earlier, $NM$ regularization is an effective tool to optimize structure. The results obtained have shown that it would yield better results. Here, we investigate the parameter-setting strategies for our $NM$ regularization.
Figure 3: The first three image shows the error of our $NM$ regularization by different parameters on the percentage of approximations(controlled by parameter $\gamma$) with $100$ nodes for the Codon Usage Dataset, the Gases Concentration Dataset and the Postural Transitions Dataset. The fourth image shows the running time of our $NM$ structure compared with DEQ on all these data sets.
The first image in Figure 3 illustrates the impact of different parameters on the error rates of our $NM$ regulation across various percentages of approximations(controlled by parameter $\gamma$) for the Codon Usage Dataset. We found that the performance can be optimized by fine-tuning the $NM$ regularization and the percentages of approximations(controlled by parameter $\gamma$). Higher approximation percentages and regularization parameters produce more Neural Modules, which help reduce overfitting problems and in turn, enhance performance as well as robustness. However, overly sparse NMs could introduce extra bias and degrade performance. After standardization, the $NM$ regularization parameter with 100 nodes performs best at $0.2$ and the approximation on top $0.05$ edges by controlling parameter $\gamma$ for the Codon Usage Dataset. However, more stable performance appears on the approximation of top $0.03$ edges.
For the $2$nd and the $3$rd image in Figure 3, we also offer this on the Gases Concentration Dataset and the Postural Transitions Dataset. We reached a similar conclusion that performance can be well optimized by tuning both the $NM$ regularization and the parameters of approximation. We observe that for these two datasets, the higher parameter $\gamma$ and $NM$ regularization parameters result in better performance. $NM$ regularization helps break large neural modules into smaller, more manageable neural modules. This segmentation leads to improved performance as previously introduced. However. compared with the Codeon Usage Data, the approximation can not to too large. On average, for 100 nodes, the $NM$ regularization parameter usually performs best at $0.2$ with the approximation on top $0.01$ edges by controlling parameter $\gamma$.
Furthermore, the fourth image in Figure 3 displays the running time of our model compared with DEQ on different complexity described by node numbers. We observed that the efficiency of $NM$ regularization is much better than DEQ. Especially for a larger number of nodes, the running time of $NM$ regularization performs well. And if we operate these Neural Modules in parallel. we would get even better results, which will be shown in the next section.
@@ 4.3 The Effect of $NM$ Regularization
To analyze the effect of $NM$ regularization, consider the probability of the edge’s weight being lower than $\gamma$ denoted as $\theta$ in the current iteration. For a neural module with $q$ nodes, the probability of a new edge integrating into the neural module would be denoted as $1-\theta^{q}$. In the case of a small $q$ value, $NM$ regularization seeks to increase the probability. Conversely, for larger $q$ values, NM regularization works to decrease the probability. Thus, NM regularization endeavors to minimize the disparity in the probability of a new edge integrating into the neural modules with varying sizes. This effect is illustrated in the first image of Figure 4.
Figure 4: The first image shows the analysis of $NM$ Regularization and the second image displays the time-consumption for different sizes of neural modules which indicates parallel computing’s advantage.
The first image in Figure 4 illustrates the analysis of $NM$ Regularization and how balanced Neural Modules are formed through this process. The second one shows the time consumption associated with different sizes of neural modules. As previously mentioned, the neural modules generated are independent and can be calculated in parallel. For instance, if a neural module comprises 50 nodes, the time to solve the system of equations required would be reduced to just 10ms in our experiments. Thus, $NM$ regularization enables sufficient management of larger node counts.
Figure 5: Neural modules generated by our $NM$ regularization for the Codon Usage Dataset with each black square denoting a neural module. The left image depicts the neural modules at the first iteration, and the right image shows the neural modules at the 10000th iteration. $NM$ regularization helps break large neural modules into smaller, more manageable neural modules.
Figure 6: Neural modules generated by our $NM$ regularization for the Gases Concentration Dataset with each black square denoting a neural module. The left image depicts the neural modules at the first iteration, and the right image shows the neural modules at the 10000th iteration. $NM$ regularization helps break large neural modules into smaller, more manageable neural modules.
Figure 7: Neural modules generated by our $NM$ regularization for the Postural Transitions Dataset with each black square denoting a neural module. The left image depicts the neural modules at the first iteration, and the right image shows the neural modules at the 10000th iteration.$NM$ regularization helps break large neural modules into smaller, more manageable neural modules.
Figure 5, Figure 6, and Figure 7 present examples of neural modules generated by our $NM$ regularization, where each black square denotes a neural module. The image on the left shows the neural modules generated at the first iteration, where the effects of $NM$ regulation are not yet apparent. The image on the right, at the 10000th iteration, illustrates how $NM$ regulation achieves independence and balance within the generated neural modules. $NM$ regularization helps break large neural modules into smaller, more manageable neural modules. This segmentation leads to improved performance and efficiency, as previously introduced.
@@ Appendix B The Connection with NNs
Variables $\mathcal{T}=\left(\begin{matrix}\mathcal{X}^{0}=[1,X^{0}]\\ \mathcal{W}^{0}=[bia_{1},W^{1}]\\ \mathcal{X}^{1}=[1,X^{1}]\\ ...\\ \mathcal{W}^{m}=[bia_{m}.W^{m}]\\ X^{m}\end{matrix}\right)$ $NM=\left(\begin{matrix}X^{0}\\ W^{1}\\ W\\ X\\ W^{m}\\ X^{m}\end{matrix}\right)$
@@
BP $\begin{cases}\nabla X^{m}\cdot g^{-1}(X^{m})\cdot W^{m}=\nabla{X}^{m-1}\\ ...\\ \nabla X^{2}\cdot f^{-1}(X^{2})\cdot W^{2}=\nabla{X}^{1}\\ \end{cases}$$\begin{cases}\sum\limits_{j\neq 1}\nabla x_{j}\cdot f^{-1}(x_{j})\cdot w_{1,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{-1}(x^{m}_{j})\cdot w^{m}_{1,j}=\nabla x_{1}\\ \sum\limits_{j\neq 2}\nabla x_{j}\cdot f^{-1}(x_{j})\cdot w_{2,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{-1}(x^{m}_{j})\cdot w^{m}_{2,j}=\nabla x_{2}\\ ...\\ \sum\limits_{j\neq m}\nabla x_{j}\cdot f^{-1}(x_{j})\cdot w_{p,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{-1}(x^{m}_{j})\cdot w^{m}_{p,j}=\nabla x_{p}\\ \end{cases}$
@@
From the table, we can see that the traditional NNs structure represents a special case of our framework where the system of equations is solved asynchronously. In the tree-like structure, the coefficient matrix of the equations is composed of the parameter matrices for each level, and these coefficients are positioned near the diagonal. In contrast, our framework generalizes the coefficient to form a full matrix across the $1$ to the $p-1$ level, indicating a transition from a tree-based structure to a general graph structure.
Appendix C The Connection with DEQ

### v3

@@ 2 Related Works
To progress the existing tree-like structure for NNs. Yuan 2020ECCV recently provided a topological perspective, highlighting the benefits of dense connections offered through shortcuts in optimization [10] [12] [15] [16] [17] [28]. Furthermore, sparsity constraints have also been proven effective in optimizing learned structures across various applications [24] [26] [33]. In their approach, the structure of NNs is organized as a DAG, whereas we organize it as a more general graph structure.
@@
The fixed point of the implicitly hidden layer can also be served as a solution [4] [5] [6] [8] [13] [31], as demonstrated in the subsequent works [20] [21] [22]. Departing from the infinite structure of implicitly hidden layers [29] [34] [35] [36] [40], we organize it into a general graph structure. Furthermore, for larger NNs, we introduce a novel regularization method to organize neural units into multiple independent neural modules. Compared with implicitly hidden layers, our method can improve the efficiency by parallel computing as well as the performance by reducing overfitting. Furthermore, it enhances the interpretability of implicitly hidden layers.
@@
Our process also involves compressing NNs. In recent years, various algorithms have been developed, including quantization [27], low-rank approximation [2] [7] [32], knowledge distillation, and network pruning [19] [23] [30] [37], etc. The network pruning method we use in this work is weight pruning, which aims to eliminate weak connections. In addition, the method also involves filter pruning that removes entire redundant filters suc9. We aim to improve weight pruning for our framework using a method similar, which evaluates the gradient at the pruned model and applies parameters’ updates to the dense model. In our framework, this process coordinates with an elegant regularization to automatically allocate Neural Modules.
@@
OptNet integrates optimization quadratic problems for nodes within the same layer [3]. However, this approach introduces bias and lacks an interpretable structure, affecting overall interpretability. Furthermore, OptNet uses parameters derived from quadratic problems for backpropagation. These parameters, determined by nodes from the previous layer, are difficult to manage in complexity. This complexity not only increases the computational burden but also exacerbates overfitting problems. Using a differentiable function for these parameters adds further bias.
@@
Graph Neural Networks (GNNs) specially address the needs of geometric deep learning [1] [9] [38] [39]. GNNs adapt their structure to the input graph, capturing complex dependencies [11] [25]. This adaptability enables them to predict properties of geometric data, which primarily deal with graphs, our innovation involves evolving the tree-like structure into a general graph format for NNs, expanding their applicability and potential.
The flexibility of this graph structure has also been investigated in studies similar to those on Reservoir Computing [41]. These studies have utilized a recurrent neural network framework where neuron connections are established randomly, where the weights remain static post-initialization. Reservoir Computing capitalizes on the dynamics of a non-linear system. In contrast, our Neural Module framework is designed around a balanced system paradigm, allowing for the adaptive learning of both weights and network structure during processing. This feature enhances our framework’s versatility and effectiveness, setting it apart from Reservoir Computing with its stronger capabilities.
@@ 3.1 Rationale for Introducing the Structure
Consider the traditional tree-like structure for an NN with $m$ layers. Let $N^{0}$ be the nodes of the first layer and the input values feed into the first layer. Let $N^{m}$ be nodes for the last layer and they feed for the output values. For any node in the other layers, $n^{k}_{i}\in N^{k},0<k<m$, it represents the intermediate values. The parameters of the model are represented as the edges in the structure. Let $e^{k}_{i,j}\in E^{k},1\leq k\leq m$ be the the edge connecting $n^{k-1}_{i}$ and $n^{k}_{j}$. Then the traditional tree-like structure for NNs which is organized as an asynchronous hierarchical structure can be formalized as follows: $\mathcal{T}=\{N^{0},E^{1},N^{1},...E^{m},N^{m}\}$.
@@ 3.3 Forward Process
Existing numerical methods such as Newton-Raphson method, can effectively solve these above equations. In real-world applications, besides Newton’s methods, the efficiency can be optimized by iterative methods,␣dichotomy, or␣secant methods. Note that, each variable is processed by an activation function $f$, making the transformation nonlinear. If we consider $\mathcal{G}$ as a Neural Module, we can introduce the following theorem. According to [14] [18], the theorem will be proved in the Appendix.
@@ 3.4 Backward Process
$\begin{cases}\sum\limits_{j\neq 1}\nabla x_{j}\cdot f^{\prime}(x_{j})\cdot w_{1,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{\prime}(x^{m}_{j})\cdot w^{m}_{1,j}=\nabla x_{1}\\ \sum\limits_{j\neq 2}\nabla x_{j}\cdot f^{\prime}(x_{j})\cdot w_{2,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{\prime}(x^{m}_{j})\cdot w^{m}_{2,j}=\nabla x_{2}\\ ...\\ \sum\limits_{j\neq m}\nabla x_{j}\cdot f^{\prime}(x_{j})\cdot w_{p,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{\prime}(x^{m}_{j})\cdot w^{m}_{p,j}=\nabla x_{p}\\ \end{cases}$
@@
$\nabla W^{T}_{:,j}=\nabla x_{j}\cdot f^{\prime}(x_{j})\cdot\mathcal{H}_{j}~{}.$ (3)
@@
$\nabla W^{m}=\nabla X^{mT}\cdot g^{\prime}(X^{mT})\cdot f(X)~{}.$ (4)
@@ 3.5 Neural Module Optimization
For the element $z_{k},1\leq k\leq p$ in $\mathcal{Z}$, initialized by the connected subgraphs formulated in the first iteration,
$z_{k}=the\ number\ of\ edges\ connected\ to\ node\ n_{k}.$ (7)
@@
Figure 2: The impact of $L1$ and $L2$ regularization is significant in neural modules. Nodes within larger neural modules have a greater number of edges, which leads to a higher value of $z$ and, consequently, a stronger regularization signal. This results in a more potent force pulling the values towards zero. The parameter $\gamma$ plays a crucial role in pruning edges with smaller weights, which not only prevents the uncontrolled growth of large neural modules but also fosters the emergence of smaller, more manageable neural modules. This is why NM (Neural Module) regularization is instrumental in creating a balanced structure within neural modules.
Based on the outcomes of $L1$ and $L2$ regularization, in the context of the $L1$ norm, the optimal value associated with the node $n_{k}$ is.
$w^{*}_{j,k}=sgn(w_{j,k})(|w_{j.k}|-\alpha\cdot z_{k})_{+}~{},$ (9)
where + takes the positive part of the value. For the $L2$ norm case, the best value associated with the node $n_{k}$ is
$w^{*}_{j,k}=\frac{1}{1+\alpha\cdot z_{k}}\cdot w_{j,k}~{}.$ (10)
As derived from Equations 9 and 10, an increased value of $z$ leads to the reduction of weights, thereby preventing the growth of large neural modules through the application of the threshold parameter $\gamma$, as previously mentioned. Concurrently, this mechanism fosters the emergence of smaller neural modules by utilizing lower values of $z$. Consequently, NM regularization contributes to the formation of well-balanced, independent neural modules, which, as discussed earlier, offer advantages for parallel computing.
@@ 40 Nodes 60 Nodes 80 Nodes 100 Nodes 200 Nodes 300 Nodes
NML&NM 0.0709 0.2416 0.0724 0.0570 0.0631 0.0635
4.2 Efficiency of Neural Modules
In this section, we delve into the optimization of Neural Module efficiency through the application of $NM$ regularization. As previously mentioned, every structure considered is a subgraph of a fully connected graph, with the initial general graph acting as the search space for our model. Our $NM$ regularization serves as a potent mechanism for structural optimization, enhancing the effectiveness and balance of neural modules. The performance of these modules, particularly for nodes numbering less than 300, was demonstrated in the preceding section, where it was shown to yield superior outcomes.
To assess the efficiency of our $NM$ regularization, we present a comparison of the model’s running time with DEQ across varying complexities, represented by node counts below 300. Figure 4, Image A, illustrates that the efficiency of NM regularization significantly outperforms DEQ, especially with a larger number of nodes. This superiority is attributed to $NM$ regularization’s ability to create multiple independent neural modules, which efficiently mitigate computational complexity.
For networks with a larger number of nodes, we can leverage the parallel processing capabilities of Neural Modules to enhance the efficiency of our framework, as previously discussed. $NM$ regularization facilitates the creation of multiple independent and well-balanced neural modules, which are inherently suited for parallel computing, particularly when transitioning to GPU-based computation. In this extension, we increase the node count from 300 to 3000 and incorporate GPU acceleration to compute the algorithms more efficiently.
[b] Refer to caption
(a) Nodes’ number up to 300. Neural modules vs.DEQ.
(b) Nodes’ number up to 3000. Neural modules parallel vs. Neural modules.
Figure 4: Neural Modules offer significant advantages over DEQ in terms of computational efficiency. Moreover, through parallel computation, $NM$ regularization provides an opportunity to extend the applicability of this framework to larger-scale applications. By harnessing the power of parallel processing, $NM$ regularization allows for the efficient handling of increased complexity, making it a promising approach for tackling more extensive computational challenges.
In Figure 4, Image B, we conducted a comparison between the running times of Neural Modules operating in parallel and Neural Modules without parallelization on large realistic models. For these experiments, we utilized 12 threads. Our results indicate that the parallel implementation of $NM$ regularization offers a computational speedup of approximately 6 to 7 times. This demonstrates the substantial efficiency gains achievable through parallel processing in the context of NM regularization.
Furthermore, for networks with a higher number of nodes, parallel $NM$ regularization also yields increasingly better results, as demonstrated in Figure 5. In our analysis, we compared methods such as DEQ, DAG, and OPTNET, and observed that when the number of nodes exceeds 300, these methods generally result in overtime. We established the average performance of these three models with node counts within 300 as our baseline for comparison.
Additionally, to assess the performance of our framework on applications with larger node numbers, we included the NN structure as a reference in our experiments. We discovered that our framework performs better on larger models. This indicates that our framework is more efficient and accurate, particularly when dealing with complex, large-scale neural network structures.
(a) Codon Usage Dataset
(b) Gases Concentration Dataset
(c) Postural Transitions Dataset
Figure 5: The performance on the 3 Datasets with larger nodes’ number.
@@ 4.3 The Effect of $NM$ Regularization
(a) Codon Usage Dataset
(b) Gases Concentration Dataset
(c) Postural Transitions Dataset
Figure 6: Neural modules generated by our $NM$ regularization for the Codon Usage Dataset, the Gases Concentration Dataset, and the Postural Transitions Transitions Dataset with each black square denoting a neural module. The left image depicts the neural modules at the first iteration, and the right image shows the neural modules at the 10000th iteration.$NM$ regularization helps break large neural modules into smaller, more manageable neural modules.
At last, Figure 6 presents examples of neural modules generated by our $NM$ regularization, where each black square denotes a neural module. The image on the left shows the neural modules generated at the first iteration, where the effects of $NM$ regulation are not yet apparent. The image on the right, at the 10000th iteration, illustrates how $NM$ regulation achieves independence and balance within the generated neural modules. $NM$ regularization helps break large neural modules into smaller, more manageable neural modules. This segmentation leads to improved performance and efficiency, as previously introduced.
@@ Appendix B The Connection with NNs
Variables $\mathcal{T}=\left(\begin{matrix}\mathcal{X}^{0}=[1,X^{0}]\\ \mathcal{W}^{0}=[bias_{1},W^{1}]\\ \mathcal{X}^{1}=[1,X^{1}]\\ ...\\ \mathcal{W}^{m}=[bias_{m}.W^{m}]\\ X^{m}\end{matrix}\right)$ $NM=\left(\begin{matrix}X^{0}\\ W^{1}\\ W\\ X\\ W^{m}\\ X^{m}\end{matrix}\right)$
@@
BP $\begin{cases}\nabla X^{m}\cdot g^{\prime}(X^{m})\cdot W^{m}=\nabla{X}^{m-1}\\ ...\\ \nabla X^{2}\cdot f^{\prime}(X^{2})\cdot W^{2}=\nabla{X}^{1}\\ \end{cases}$$\begin{cases}\sum\limits_{j\neq 1}\nabla x_{j}\cdot f^{\prime}(x_{j})\cdot w_{1,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{\prime}(x^{m}_{j})\cdot w^{m}_{1,j}=\nabla x_{1}\\ \sum\limits_{j\neq 2}\nabla x_{j}\cdot f^{\prime}(x_{j})\cdot w_{2,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{\prime}(x^{m}_{j})\cdot w^{m}_{2,j}=\nabla x_{2}\\ ...\\ \sum\limits_{j\neq m}\nabla x_{j}\cdot f^{\prime}(x_{j})\cdot w_{p,j}+\sum\limits_{j=1}^{|N^{m}|}\nabla x^{m}_{j}\cdot g^{\prime}(x^{m}_{j})\cdot w^{m}_{p,j}=\nabla x_{p}\\ \end{cases}$
@@
From the table, we can see that the traditional NN structure represents a special case of our framework where the system of equations is solved asynchronously. In the tree-like structure, the coefficient matrix of the equations is composed of the parameter matrices for each level, and these coefficients are positioned near the diagonal. In contrast, our framework generalizes the coefficient to form a full matrix across the $1$ to the $p-1$ level, indicating a transition from a tree-based structure to a general graph structure.
Appendix C The Motivation of Reorganizing the Neural Networks from Tree-like Structure to General Graph Structure
For the traditional tree-like structure, the asynchronous forward and backward process can also be considered to solve a system of equations as shown in Table 4. In our framework, the feed-forward processes for traditional tree-like structures are essential to solve the equations as shown in Figure 9, which is similar to the backward process.
For traditional tree-like structures, the asynchronous forward and backward processes can indeed be utilized to address the solution of a system of equations, as illustrated in Table 4. In our framework, the feed-forward processes for traditional tree-like structures are essential to solve the equations as shown in Figure 9, This process bears resemblance to the backward propagation phase.
$\begin{cases}w^{1}_{1}+\sum\limits_{j=1}^{|N^{0}|}x^{0}_{j}\cdot w^{1}_{j,1}=x^{1}_{1}\\ w^{1}_{2}+\sum\limits_{j=1}^{|N^{0}|}x^{0}_{j}\cdot w^{1}_{j,2}=x^{1}_{2}\\ ...\\ w^{1}_{|N^{1}|}+\sum\limits_{j=1}^{|N^{0}|}x^{0}_{j}\cdot w^{1}_{j,|N^{1}|}=x^{1}_{|N^{1}|}\\ ......\\ w^{m-1}_{1}+\sum\limits_{j=1}^{|N^{m-2}|}f(x^{m-2}_{j})\cdot w^{m-1}_{j,1}=x^{m-1}_{1}\\ w^{m-1}_{2}+\sum\limits_{j=1}^{|N^{m-2}|}f(x^{m-2}_{j})\cdot w^{m-1}_{j,2}=x^{m-1}_{2}\\ ...\\ w^{m-1}_{|N^{m-1}|}+\sum\limits_{j=1}^{|N^{m-2}|}f(x^{m-2}_{j})\cdot w^{m-1}_{j,|N^{m-1}|}=x^{m-1}_{|N^{m-1}|}\\ \end{cases}$
Figure 9: Tadition tree-like structure for NN is essentially to solve these equations in our framework.
Assign the input and bias-related values to the right side. Allocate the values associated with each node to the left side. Consequently, the equations are reformulated as shown in Figure 10.
$\begin{cases}-x^{1}_{1}=-w^{1}_{1}-\sum\limits_{j=1}^{|N^{0}|}x^{0}_{j}\cdot w^{1}_{j,1}\\ -x^{1}_{2}=-w^{1}_{2}-\sum\limits_{j=1}^{|N^{0}|}x^{0}_{j}\cdot w^{1}_{j,2}\\ ...\\ -x^{1}_{|N^{1}|}=-w^{1}_{|N^{1}|}-\sum\limits_{j=1}^{|N^{0}|}x^{0}_{j}\cdot w^{1}_{j,|N^{1}|}\\ ......\\ \sum\limits_{j=1}^{|N^{m-2}|}f(x^{m-2}_{j})\cdot w^{m-1}_{j,1}-x^{m-1}_{1}=-w^{m-1}_{1}\\ \sum\limits_{j=1}^{|N^{m-2}|}f(x^{m-2}_{j})\cdot w^{m-1}_{j,2}-x^{m-1}_{2}=-w^{m-1}_{2}\\ ...\\ \sum\limits_{j=1}^{|N^{m-2}|}f(x^{m-2}_{j})\cdot w^{m-1}_{j,|N^{m-1}|}-x^{m-1}_{|N^{m-1}|}=w^{m-1}_{|N^{m-1}|}\\ \end{cases}$
Figure 10: Assign the values associated with inputs and bias to the right-hand side. Place the values related to each node on the left-hand side. This arrangement allows us to extract the coefficient $\mathcal{C}$ for our framework.
Then, the coefficient $\mathcal{C}$ for the tree-like structure in our framework would be as Image A in Figure 11.
$\begin{pmatrix}-E^{1}&0&0&0&...&0&0\\ W^{1}&-E^{2}&0&0&...&0&0\\ 0&W^{2}&-E^{3}&0&...&0&0\\ 0&0&W^{3}&-E^{4}&...&0&0\\ ......\\ 0&0&0&0&...&W^{m-2}&E^{m-1}\\ \end{pmatrix}$
(a) The Coefficient $\mathcal{C}$ for the Tree-like Structure and $E$ is the identity matrix for each level.
$\begin{pmatrix}-E^{1}&0&0&0&...&0&0\\ W^{1}&-E^{2}&0&0&...&0&0\\ V^{1}&W^{2}&-E^{3}&0&...&0&0\\ V^{1}&V^{2}&W^{3}&-E^{4}&...&0&0\\ ......\\ V^{1}&V^{2}&V^{3}&V^{4}&...&W^{m-2}&E^{m-1}\\ \end{pmatrix}$
(b) The Coefficient $\mathcal{C}$ for the DAG or ResNet Structure and $E$ is the identity matrix for each level.
Figure 11: The Coefficient $\mathcal{C}$ for traditional NN Structure
The tree-like structure in our framework is indeed more constrained for each node within the structure. According to the Universal Approximation Theorem, the flexibility of a Neural Network (NN) is contingent upon the number of neurons within a single hidden layer. This implies that for any nodes in a tree structure if they are solely dependent on the neurons in the preceding layer, their capacity for complex function approximation is severely limited. Typically, nodes in the first layer are represented as a linear transformation of the input values.
Some existing work has attempted to generalize the structure to a Directed Acyclic Graph (DAG), such as in the case of ResNet. In these architectures, additional weights are introduced to the Lower Triangular of the coefficient matrix $\mathcal{C}$. Let these weights be represented by $V$. Consequently, the matrix $\mathcal{C}$ is depicted in Image B of Figure 11. This modification allows for a more flexible and expressive model, which can better approximate complex functions and handle larger datasets, aligning with the theorem’s assertion that a single hidden layer with a sufficient number of neurons can approximate any continuous function on a compact subset of the real numbers.
The Lower Triangular Coefficient matrix indeed reaches the limits of asynchronous structures. However, the potential of the Upper Triangular matrix remains largely untapped. In our work, we extend the asynchronous structure to a synchronous one, breaking through the limitations of the Lower Triangular and generalizing the coefficient matrix $\mathcal{C}$. Within our Neural Module framework, the $\mathcal{C}$ is depicted in Figure 12, showcasing a more comprehensive and interconnected structure that allows for greater flexibility and performance.
$\begin{pmatrix}-1&w_{2,1}&w_{3,1}&w_{4,1}&...&w_{p-1,1}&w_{p,1}\\ w_{1,2}&-1&w_{3,2}&w_{4,2}&...&w_{p-1,2}&w_{p,2}\\ w_{1,3}&w_{2,3}&-1&w_{4,3}&...&w_{p-1,3}&w_{p,3}\\ w_{1,4}&w_{2,4}&w_{3,4}&-1&...&w_{p-1,4}&w_{p,4}\\ ......\\ w_{1,p}&w_{2,p}&w_{3,p}&w_{4,p}&...&w_{p-1,p}&-1\\ \end{pmatrix}$
Figure 12: The Coefficient $\mathcal{C}$ for our NM Structure
In our framework, we have enhanced the representational capacity of each neuron, thereby unlocking the full potential of Neural Networks (NNs). Concurrently, we have eliminated the structural bias that is typically inherent in predefined structures such as traditional tree structures or Directed Acyclic Graphs (DAGs). This innovation allows our framework to be more adaptable and less constrained by the limitations of fixed architectural biases, leading to a more flexible and effective NN design.
For large coefficient matrices $\mathcal{C}$, solving the system of equations can indeed be challenging. Our framework addresses this by proposing $NM$ regularization, which serves as an approximation method for $\mathcal{C}$ and incorporates parallel computation to enhance efficiency. However, it’s important to note that $NM$ regularization is an approximation technique, and it operates on the complete set of the coefficient matrix $\mathcal{C}$ as previously introduced. This approach allows us to break down the complex system into more manageable parts, which can then be solved in parallel, significantly reducing the computational burden and improving the overall performance of the framework.
Appendix D The Connection with DEQ